### v1

@@ Abstract
Digital health interventions are increasingly used in physical and occupational therapy to deliver home exercise programs via sensor-equipped devices such as smartphones, enabling remote monitoring of adherence and performance. However, in the current digital intervention paradigm, treatments in the form of software are typically programmed before clinical encounters by software engineers as libraries of parametrized exercise modules targeting broad impairment categories. Consequently, at the point of care, clinicians can only select modules and adjust a narrow set of parameters such as duration and repetitions. Patient-specific needs that emerge during encounters—such as distinct movement limitations, personal goals, or home and work constraints—are rarely reflected in the software, limiting personalization and contributing to lower adherence and reduced therapeutic benefit. In this study, we propose a new digital intervention paradigm that uses large language models (LLMs) to translate clinicians’ exercise prescriptions into intervention software. In a prospective single-arm feasibility study with 20 licensed physical and occupational therapists and a standardized patient, clinicians created 40 individualized upper extremity programs (398 instructions) that were automatically translated into executable software. We used a standardized patient because the safety of LLM-generated intervention software has not previously been studied. Our results show that the proposed paradigm yields a 45% increase in the proportion of personalized prescriptions that can be implemented as software compared with the status quo, with unanimous consensus among therapists on its ease of use. The LLM-generated software correctly delivered 99.78% (397/398) of instructions as prescribed and monitored performance with 88.4% (352/398) accuracy; 90% (18/20) of therapists judged it safe to interact with patients, and 75% (15/20) expressed willingness to adopt it in their clinical practices. To our knowledge, this is the first prospective evaluation of clinician-directed intervention software generation with an LLM in healthcare, demonstrating feasibility and motivating larger trials to assess clinical effectiveness and safety in real-patient populations.
@@ 1 Introduction
Digital health interventions (DHIs) [Wienert2022, Tan2024] in physical and occupational therapy increasingly take the form of clinician-prescribed intervention software deployed to a patient’s smartphone or connected wearable device [Zangger2023, Krohn2024]. Such software delivers step-by-step exercise instructions and, critically, can objectively capture adherence and performance using on-device sensors (e.g., camera-based motion tracking on phones or inertial sensing in wearables). As a result, it can quantify whether exercises were completed and how they were performed—such as range of motion, repetition count, tempo, and posture—and return concise summaries to clinicians for timely adjustment of therapy (e.g., modifying difficulty or introducing new exercises). This capability represents a practical advance over standard paper exercise worksheets, which cannot capture how patients perform prescribed exercises at home between visits.
Despite these gains, the prevailing digital prescription paradigm with monitoring remains rigid in its capacity to personalize therapy. In current DHI platforms, prior to clinical encounters, therapeutic exercises are encoded as a library of parametrized intervention software, targeting broad populations with common impairments. Consequently, patient-specific needs identifiable only during encounters (for example, distinct deficits, personal goals, or home and work environments) are not fully reflected in the software. Therapists can personalize therapy in limited ways by selecting the most relevant subset of software from the provided library and adjusting a small set of parameters, such as difficulty [gauthier2022vigorous, prieto2024activehip], repetitions [cui2023lowback, malliaras2020intel], and duration [weise2022backpain]. This paradigm makes it difficult to flexibly modify exercise instructions—for example, to introduce contingencies when a patient cannot perform a prescribed motion or to incorporate patient-specific household objects into exercises. Prior studies have associated limited personalization in DHIs with lower patient adherence and reduced therapeutic effectiveness [Palazzo2016, Mahmood2023, Ricke2023]. At the same time, merely expanding configurable options to enhance personalization can impose excessive cognitive load on clinicians in routine care [Borges2023, vanTilburg2024, Pearce2024].
@@
We posit that advances in large language models (LLMs), particularly their ability to generate software from natural language instructions [llm_code_gen], can offer a viable solution to this limitation. While LLMs have already been used in several healthcare settings [busch2025llmreview, alkalbani2025llmreview], their use for generating patient-facing software remains nascent. Early work has applied LLM-based software generation to data retrieval [Shi2024], data processing [Ahn2024], and visualization [Liu2025], but not to the direct creation of clinician-prescribed intervention software. To address the constraints of current DHIs, we propose a digital prescription paradigm that leverages an LLM as an intervention software generation tool. In this paradigm, clinicians design exercises during clinical encounters, tailoring them to each patient’s physical impairments, rehabilitation goals, and home environment, and the LLM translates these prescriptions into intervention software. By freeing clinicians from pre-defined software templates, this approach allows the generated software to be structurally aligned with individualized patient needs. The resulting intervention software is then deployed to patients’ devices, where it delivers step-by-step instructions, monitors completion of each instruction, and transmits adherence and performance reports back to clinicians.
@@
Here, we report a prospective, single-arm study involving 20 licensed physical and occupational therapists and a standardized patient to assess the feasibility of clinician-directed, LLM-based software generation from clinicians’ perspectives. We employed a standardized patient because the safety of LLM-generated intervention software had not previously been investigated, and we sought an initial feasibility and safety assessment before enrolling real patients. To our knowledge, the prospective feasibility of clinician-directed LLM software generation has not been evaluated previously in any domain of healthcare. In our study (Fig. 1), each therapist conducted a therapy session with a standardized patient presenting with upper extremity weakness, during which they designed individualized upper extremity exercises comprising 40 exercise programs and 398 total instructions. These exercises were immediately translated into intervention software by the LLM. Following the session, therapists observed the software instruct the patient, monitor completion of each instruction, and report monitored results back to them for evaluation. Therapists evaluated the LLM-based digital prescription paradigm across five dimensions: (i) usability of intervention software generation via the LLM, (ii) flexibility in accommodating personalized exercise design compared with the existing digital intervention paradigm, (iii) generated software’s accuracy of instruction delivery and monitoring, (iv) perceived safety in patient use, and (v) clinician acceptance of LLM-based software generation.
@@ 2.2 Flexibility in Accommodating Personalized Interventions
Clinician-generated intervention software delivered 45% more personalized interventions than the status quo paradigm.
In the status quo digital intervention paradigm, physical and occupational therapists prescribe exercises by selecting from a library of intervention software that has been programmed prior to the clinical encounter. By contrast, we propose a paradigm in which therapists generate intervention software after the encounter by writing free-form, patient-specific exercise descriptions that are then compiled into executable programs (Fig.1). We hypothesized that allowing clinicians to directly generate intervention software, informed by their current understanding of the patient’s needs, would enable a larger fraction of personalized prescriptions to be represented in software than the status quo template-based approach. To validate this hypothesis, we analyzed whether each personalized exercise prescribed after the standardized patient session could be translated into intervention software under (i) our proposed LLM-based paradigm and (ii) existing digital rehabilitation interventions (Sec.4.6.2).
Our proposed paradigm successfully translated 100% (40/40) of personalized exercise prescriptions into intervention software, whereas the existing paradigm could represent only 55% (22/40). This corresponds to an absolute increase of 45 percentage points in the proportion of personalized prescriptions that can be delivered as software under our paradigm compared with the status quo. Table 1 summarizes the personalization features that could not be encoded within the existing paradigm (the analysis data are provided in Supplement E). The most frequent limitation was procedural variation, where therapists substantially modified exercise instruction sequence, movement quality, or goals (37.5%, 15/40; for example, replacing simple shoulder abduction–adduction with a functional cube-transfer task). Other common limitations included intensity customization via new equipment (15%, 6/40; for example, adding a rubber band to increase resistance during finger abduction–adduction) and novel contingency logic beyond predefined progressions (15%, 6/40; for example, progressing difficulty only when new criteria were met rather than based on fixed repetitions). Finally, several prescriptions specified compensatory strategies that the template could not represent (10.0%, 4/40; for example, instructing the patient to rotate the trunk or use shoulder adduction to facilitate task performance).
Together, these findings suggest that clinician-driven software generation substantially expands the expressive capacity of digital interventions, enabling a larger share of clinically meaningful personalization to be preserved when translating therapists’ prescriptions into intervention software.
@@ 2.3 Instruction and Monitoring Accuracy of Software
While LLM enables therapists to flexibly personalize the intervention software, we also observed some hallucinations in the generated software.
@@
The LLM-generated intervention software guided the prescribed exercises and monitored the completion of each instruction. Therapists then reviewed the software-generated performance report to assess its monitoring accuracy across all 398 instructions (Fig. 1(iii), and refer to Sec  4.6). The software’s monitoring accuracy, sensitivity, and specificity were 88.4% (95% CI, 0.843–0.915), 88.6% (95% CI, 0.836–0.923), and 87.5% (95% CI, 0.809–0.921), respectively. Confidence intervals were estimated using a generalized linear mixed model (GLMM) [bolker2009glmmguide] to account for the hierarchical structure of the experiment, in which each exercise goal (e.g., shoulder flexion; exerciseID) was assigned to a therapist (clinicianID), and each therapist created multiple instructions (stepID) for that goal. The GLMM incorporated potential intra-class correlations (ICCs) among exercises, therapists, and instructions when computing accuracy (proportion of correct classifications), sensitivity (proportion of actual successes correctly identified), and specificity (proportion of actual failures correctly identified). To assess the impact of ICCs, we also calculated Wilson score intervals  [brown2001intervals], which do not account for correlations. As shown in Fig. 2, the similarity of the confidence intervals between the two methods indicates that the effects of correlations were minimal. No syntax error was encountered with any LLM-generated software. The observed monitoring inaccuracies arose from two sources: hallucinated conditions in the LLM-generated software and errors in visual movement detection. Upon manual inspection, we identified hallucinations in the monitoring logic for 10 of 398 instructions, where the generated code checked for conditions that were not specified in the therapist’s prescribed exercise instructions. The remaining inaccuracies predominantly reflected limitations of the computer vision models used to interpret the patient’s movements, leading to incorrect classifications of whether exercises were performed as instructed.
@@ 2.4 Perceived Safety of Intervention Software in Patient Use
90% of therapists perceived software’s interactions with patients to be safe.
As shown in Fig. 1(iv), therapists observed the LLM-generated intervention software articulating and monitoring each instruction to assess the perceived safety of interaction between the software and patients. 90% (18/20) therapists reported that exercising with the software delivered via patients’ personal electric devices posed no greater risk of injury than using traditional paper worksheets which are the most common form of exercise prescription in standard practice (Fig. 3c). However, 10% (2/20) disagreed with this statement, citing concerns that discrepancies in pacing instructions (ranging from 1 to 20 seconds before proceeding to the next instruction) might place undue physical demand on some patients.
Pacing refers to the software’s ability to adaptively wait for patients to complete each instruction before advancing to the next, and it is inherently dependent on the monitoring accuracy of individual instructions. During the software’s interaction with patients, therapists evaluated and flagged the instructions with inadequate pacing. Therapists judged 92.8% (369/398) of instructions as appropriately paced. These inadequate pacing resulted from the aforementioned hallucinations in monitoring logic in the software and inaccuracies in the computer vision models. False negatives (i.e., the software failed to detect a completed movement) led to pacing delays, while false positives (i.e., the incomplete movement was incorrectly judged as complete) caused the software to advance prematurely.
Therapists noted that pacing discrepancies, whether too slow or too fast, could lead to unnecessary exertion. For instance, if the software delayed progression while verifying completion of an instruction such as “hold an object in the air,” patients might hold the position longer than needed, risking fatigue or discomfort.
@@ 3 Discussion
We propose a novel digital intervention paradigm in which clinicians use LLMs to translate their exercise prescriptions into intervention software that can remotely instruct and monitor patients’ exercises. In our study, the generated software was executed on a standardized patient’s device to provide step-by-step guidance, monitor performance, and return monitored outcomes to clinicians for review. Our findings demonstrate that clinician-driven intervention software generation with LLM is technically feasible and acceptable to therapists in a standardized-patient setting, suggesting potential for future clinical adoption pending evaluation in real-patient trials.
@@
The proposed digital intervention paradigm enabled a substantially larger fraction of personalized prescriptions to be delivered as software than the existing template-based paradigm, with a 45 percentage point increase in the proportion of prescribed exercises that could be implemented (100% vs. 55%). This finding is consistent with our paradigm’s design: therapists translate exercise prescriptions—step-by-step instructions for specific exercises—after clinical encounters. During these encounters in the study, therapists elicited individual needs (for example, physical deficits, home and work environments, rehabilitation goals) and crafted prescriptions accordingly. By contrast, under the status quo, intervention software is implemented prior to clinical encounters, and personalization is largely limited to selecting software and adjusting a narrow set of parameters (for example, repetitions or duration) while the overall exercise instruction remains fixed. Our paradigm enabled forms of personalization not previously supported by the generalized DHI template, such as incorporating patient-specific household objects used in activities of daily living and providing contingency instructions when patients had difficulty following particular motions.
@@
Within this paradigm, the LLM-generated intervention software achieved 99.7% accuracy in delivering exercise instructions (397/398 instructions delivered as prescribed). Importantly, therapists preserved their role as clinical decision makers. As in standard practice, they reviewed the patient chart (here, a clinical vignette summarizing physical deficits), interacted with the standardized patient to elicit needs, and prescribed personalized exercises. The LLM functioned solely as a translator and did not participate in clinical decision-making. As specified in the system prompt (Fig.1), the LLM was instructed to deliver exercise instructions verbatim as prescribed, permitting only minor edits (for example, resolving ambiguous pronouns such as “this” or “that” based on prior context). Because therapists determined the intervention design, this approach leverages the LLM’s strength in converting text description into structured logic[llm_code_gen] while limiting its role to code generation. Our use of LLMs is therefore distinct from prior applications that rely on more complex clinical reasoning—such as diagnosis [liu2025generalist], treatment planning [sandmann2025deepseek], or risk management [pais2024medic]. Therapists’ feedback further indicated that this constrained, code-generation role was easy to integrate into their existing clinical workflow and, in some cases, perceived as enhancing efficiency.
@@
While our proposed paradigm equipped clinicians with a tool to prescribe personalized intervention software, we also observed monitoring inaccuracies arising from both hallucinations in the generated software logic and limitations of the computer vision models used in this study. In existing template-based DHIs, libraries of intervention software are typically manually implemented, inspected, and tested by software engineers before deployment, which may, in principle, support higher monitoring reliability than an automatically generated program. By contrast, our current implementation did not include a manual validation step to mitigate suboptimal monitoring logic. Despite this limitation, we observed an 88.4% monitoring accuracy (352/398 instructions; 95% CI, 0.84–0.92), which we consider a reasonable initial outcome that warrants further investigation of LLM-based software generation for digital interventions. Moreover, the commercial LLM and computer vision models used in our study were not fine-tuned for this specific rehabilitation context. Fine-tuning these models on domain-specific data, incorporating more robust pose-estimation and activity-recognition pipelines, is likely to improve monitoring performance, especially as underlying LLM and computer vision capabilities continue to advance.
@@
Regarding patient safety, 90% (18/20) of therapists judged the LLM-generated intervention software to be safe in its patient interactions in this standardized-patient setting. Because the safety of LLM-generated intervention software had not previously been evaluated, we deliberately used a standardized patient rather than real patients in this first feasibility study. This initial safety signal is encouraging, but it should be interpreted as preliminary and will require confirmation in trials involving real patients and longer-term use. Therapists observed the software articulate each prescribed instruction and wait until monitoring confirmed completion or time-to-completion expired before proceeding. This monitoring enabled adaptive pacing aligned with patient performance, making sessions more interactive than a patient reading step-by-step instructions on a screen. Most therapists considered the pacing adequate; however, two raised safety concerns, noting premature or delayed instruction transitions that could pose risks—for example, delayed detection of completing a task such as “hold an object in the air” could lead to fatigue or muscle strain. Improving monitoring accuracy through targeted fine-tuning may reduce premature transitions caused by false positives, while delayed transitions may be mitigated by carefully selecting sensor data processing models—exposed to the LLM as APIs in the system prompt (Fig. 1)—with explicit consideration of processing latency.
@@
To broadly support physical and occupational therapy with the proposed LLM-based paradigm, our study found that iterative co-design of the system prompt (Fig. 1) between software developers and clinicians was essential. The prompt specifies guidelines for the LLM to generate intervention software, detailing desired functionalities and guardrails aligned with clinical needs. Because authoring such prompts currently requires programming expertise, AI researchers or software engineers remain necessary contributors to developing this class of tools, as with other LLM-based clinical systems. In this study, however, the prompt only partially met therapist expectations, tempering adoption. Some therapists sought finer control over instruction pacing, while others expected monitoring beyond binary labels (complete vs. incomplete). Although 75% (15/20) of therapists indicated willingness to adopt the LLM software-generation tool, with unanimous consensus (20/20) on its ease of use, these unmet needs contributed to hesitancy, underscoring the importance of early needs-finding and iterative prompt refinement to ensure alignment with clinical practice.
@@
Our findings may have implications for other clinical domains adopting DHIs. Template-based DHIs in mental health [personalization-ibis-review], chronic disease management [ai-diabetes-management], pain management [personalizing-pain-management], and speech–language pathology [digital-aphasia-therapy] exhibit similar limitations in personalization. As an illustrative (untested) example, a study of exposure and response therapy for obsessive–compulsive disorder (OCD) [Lenhard2017] did not permit clinicians to introduce new task-specific contingencies (for example, “proceed to Step 3 only if hands remain unwashed and visible for 10 minutes”) beyond predefined instructions. In contrast, our approach allow clinicians to encode such contingencies directly in the intervention software (Table 1). Prospective studies are needed to validate whether the benefits of LLM-based software generation observed in this study extend to other areas of healthcare.

### v2

@@ Abstract
Digital health interventions are increasingly used in physical and occupational therapy to deliver home exercise programs via sensor-equipped devices such as smartphones, enabling remote monitoring of adherence and performance. In the current digital intervention paradigm, however, exercise software is typically programmed before clinical encounters as libraries of pre-defined modules aimed at broad impairment categories. At the point of care, clinicians can only select from this library and adjust a narrow set of parameters such as duration and repetitions. Patient-specific needs that emerge during encounters—such as distinct movement limitations, personal goals, or home and work constraints—are rarely reflected in the software, limiting personalization and contributing to lower adherence and reduced therapeutic benefit. In this study, we propose a digital intervention paradigm in which large language models (LLMs) act as constrained translators that convert clinicians’ exercise prescriptions into intervention software. Clinicians preserve their role as clinical decision-makers: they design exercises during the encounter, tailored to each patient’s impairments, goals, and environment, and the LLM generates software to match these prescriptions. We conducted a prospective single-arm feasibility study with 20 licensed physical and occupational therapists and a standardized patient. Clinicians created 40 individualized upper extremity exercise programs (398 total instructions), which were automatically translated into executable software. A standardized patient was used because the safety of LLM-generated intervention software had not previously been evaluated. We benchmarked each therapist-designed prescription against a representative template-based DHI to estimate how many could be delivered under the status quo paradigm. The proposed paradigm yielded a 45% increase in the proportion of personalized prescriptions that could be implemented as software compared with the status quo (100% vs 55%; p <0.01), with unanimous agreement among therapists that it was easy to use. The LLM-generated software correctly delivered 99.7% (397/398) of instructions as prescribed and monitored performance with 88.4% (95% CI, 0.843–0.915) accuracy. Overall, 90% (18/20) of therapists judged the system safe for patient interaction, and 75% expressed willingness to adopt it in their clinical practice. To our knowledge, this is the first prospective evaluation of clinician-directed intervention software generation with an LLM in healthcare, demonstrating feasibility and motivating larger trials to assess clinical effectiveness and safety in real-patient populations.
@@ 1 Introduction
Digital health interventions (DHIs) [Wienert2022, Tan2024] in physical and occupational therapy increasingly take the form of clinician-prescribed intervention software deployed to a patient’s smartphone or connected wearable device [Zangger2023, Krohn2024]. Such software delivers step-by-step exercise instructions and can objectively capture adherence and performance using on-device sensors (e.g., camera-based motion tracking on phones or inertial sensing in wearables). As a result, it can quantify whether exercises were completed and how they were performed—such as range of motion, repetition count, tempo, and posture—and return concise summaries to clinicians for timely adjustment of therapy (e.g., modifying difficulty or introducing new exercises). This capability represents a practical advance over standard paper exercise worksheets, which cannot capture how patients perform prescribed exercises at home between visits.
Despite these gains, the prevailing digital prescription paradigm with monitoring remains rigid in its capacity to personalize therapy. In current DHI platforms, prior to clinical encounters, therapeutic exercises are encoded as a library of parametrized intervention software, targeting broad populations with common impairments. Consequently, patient-specific needs identifiable only during encounters (for example, distinct deficits, personal goals, or home and work environments) are not fully reflected in the software. Therapists can personalize therapy in limited ways by selecting the most relevant subset of software from the provided library and adjusting a small set of parameters, such as difficulty [gauthier2022vigorous, prieto2024activehip], repetitions [cui2023lowback, malliaras2020intel], and duration [weise2022backpain]. This rigid paradigm makes it difficult to flexibly modify exercise instructions—for example, to introduce contingencies when a patient cannot perform a prescribed motion or to incorporate patient-specific house layout into activities of daily living. Prior studies have associated limited personalization in DHIs with lower patient adherence and reduced therapeutic effectiveness [Palazzo2016, Mahmood2023, Ricke2023]. At the same time, simply expanding configurable options to enhance personalization can impose excessive cognitive load on clinicians in routine care [Borges2023, vanTilburg2024, Pearce2024], creating a tension between richer tailoring and practical usability.
@@
We posit that advances in large language models (LLMs), particularly their ability to generate software from natural language instructions [llm_code_gen], can offer a viable way to resolve this tension. While LLMs have already been used in several healthcare settings [busch2025llmreview, alkalbani2025llmreview], their use for generating patient-facing software remains nascent. Early work has applied LLM-based software generation to data retrieval [Shi2024], data processing [Ahn2024], and visualization [Liu2025], but not to the direct creation of clinician-prescribed intervention software. To address the constraints of current DHIs, we propose a digital prescription paradigm in which the LLM functions as a constrained translator rather than a clinical decision-maker. In this paradigm, clinicians design exercises during clinical encounters, tailoring them to each patient’s physical impairments, rehabilitation goals, and home environment. Then, the LLM converts these prescriptions into intervention software without altering their clinical content. By freeing clinicians from pre-defined software templates—while preserving their role as the sole authors of exercise prescriptions—this approach inverts the status quo sequence, generating software after the clinician has full knowledge of the patient and allowing the resulting programs to be structurally aligned with individualized patient needs.
@@
We report a prospective, single-arm study involving 20 licensed physical and occupational therapists and a standardized patient to assess the feasibility of clinician-directed, LLM-based software generation from clinicians’ perspectives. We employed a standardized patient because the safety of LLM-generated intervention software had not previously been investigated, and we sought an initial feasibility and safety assessment before enrolling real patients. To our knowledge, the prospective feasibility of clinician-directed LLM software generation has not been evaluated previously in any domain of healthcare. In our study (Fig. 1), each therapist conducted a therapy session with a standardized patient, during which they designed individualized upper extremity exercises comprising 40 exercise programs and 398 total instructions. These exercises were immediately translated into intervention software by the LLM. We then benchmarked the resulting prescriptions against a representative template-based DHI to estimate how many could be delivered under the status quo paradigm and to compare the personalization supported by each approach. Following the session, therapists observed the software instruct the patient, monitor completion of each instruction, and report monitored results back to them for evaluation. Therapists evaluated the LLM-based digital prescription paradigm across five dimensions: (i) usability of intervention software generation via the LLM, (ii) flexibility in accommodating personalized exercise design compared with the existing digital intervention paradigm, (iii) generated software’s accuracy of instruction delivery and monitoring, (iv) perceived safety in patient use, and (v) clinician acceptance of LLM-based software generation.
@@ 2.2 Flexibility in Accommodating Personalized Interventions
Proposed paradigm implemented 45% more personalized prescriptions than current DHIs.
In the current digital intervention paradigm, physical and occupational therapists personalize care by retrofitting their prescriptions into a pre-programmed library of intervention software written before the clinical encounter. They choose the closest available program and adjust a few parameters (for example, repetitions or difficulty), but cannot change the underlying instructions. In our paradigm, therapists instead write free-text, patient-specific exercise prescriptions after the encounter, and an LLM converts these prescriptions into executable software (Fig.1). We therefore hypothesized that this post-encounter, clinician-directed software generation would allow a larger fraction of personalized prescriptions to be implemented as software than the library-based, retrofit approach used in existing DHIs. To test this, we examined whether each personalized exercise prescribed after the standardized-patient session could be translated into intervention software under (i) our LLM-based paradigm and (ii) a generalized representation of current DHIs (Sec. 4.6.2).
Under our proposed paradigm, all 40 personalized prescriptions (100%) could be translated into intervention software, compared with 22/40 (55%) under the generalized existing DHI template (Fisher’s exact test, $p<0.01$). All LLM-generated intervention software compiled without syntax errors and were executable. Table 1 summarizes the personalization features that could not be encoded within the existing paradigm (analysis data in Appendix E). The most frequent limitation was procedural variation, where therapists substantially changed the exercise procedure—for example, altering the sequence, movement quality, or goals, such as replacing simple shoulder abduction–adduction with a functional cube-transfer task (37.5%, 15/40). Other common limitations included incorporating new household objects not represented in the template (15%, 6/40; for example, adding a rubber band to increase resistance during finger abduction–adduction) and contingency rules to cope with difficulties in movement due to patient-specific deficits (15%, 6/40). Several prescriptions also specified compensatory strategies the template could not encode (10.0%, 4/40; for example, instructing the patient to rotate the trunk or use shoulder adduction to facilitate task performance).
Together, these findings suggest that clinician-directed software generation substantially expands what digital interventions can express, preserving a larger share of clinically meaningful personalization when translating therapists’ prescriptions into intervention software.
@@ 2.3 Instruction and Monitoring Accuracy of Software
The LLM-generated intervention software guided the prescribed exercises and monitored the completion of each instruction. Therapists then reviewed the software-generated performance report to assess its monitoring accuracy across all 398 instructions (Fig. 1(iii), and refer to Sec. 4.6.1). The software’s monitoring accuracy, sensitivity, and specificity were 88.4% (95% CI, 0.843–0.915), 88.6% (95% CI, 0.836–0.923), and 87.5% (95% CI, 0.809–0.921), respectively. Confidence intervals were estimated using a generalized linear mixed model (GLMM) [bolker2009glmmguide] to account for the hierarchical structure of the experiment, in which each exercise goal (e.g., shoulder flexion; exerciseID) was assigned to a therapist (clinicianID), and each therapist created multiple instructions (stepID) for that goal. The GLMM incorporated potential intra-class correlations (ICCs) among exercises, therapists, and instructions when computing accuracy (proportion of correct classifications), sensitivity (proportion of actual successes correctly identified), and specificity (proportion of actual failures correctly identified). To assess the impact of ICCs, we also calculated Wilson score intervals  [brown2001intervals], which do not account for correlations. As shown in Fig. 2, the similarity of the confidence intervals between the two methods indicates that the effects of correlations were minimal. The observed monitoring inaccuracies arose from two sources: hallucinated conditions in the LLM-generated software and errors in visual movement detection. Upon manual inspection, we identified hallucinations in the monitoring logic for 2.5% (10/398) of instructions, where the generated code checked for conditions that were not specified in the therapist’s prescribed exercise instructions. The remaining inaccuracies predominantly reflected limitations of the computer vision models used to interpret the patient’s movements, leading to incorrect classifications of whether exercises were performed as instructed.
@@ 2.4 Perceived Safety of Intervention Software in Patient Use
90% of therapists perceived software–patient interactions to be safe.
As shown in Fig. 1(iv), therapists observed the LLM-generated intervention software delivering and monitoring each instruction to assess the safety of its interactions with patients. Overall, 90% (18/20) of therapists reported that exercising with the software on patients’ personal electronic devices posed no greater risk of injury than using traditional paper worksheets, which remain the most common format for home exercise prescription in standard practice (Fig. 3c). The remaining 10% (2/20) disagreed, primarily due to concerns about pacing discrepancies—instances in which the software advanced too quickly or too slowly between instructions.
In this context, pacing refers to the software’s ability to wait for patients to complete each instruction before progressing, based on its monitoring of movement. During the interaction, therapists evaluated pacing for each instruction and flagged those they judged inadequate. Overall, 92.8% (369/398) of instructions were rated as appropriately paced. Inadequate pacing arose from hallucinations in the generated monitoring logic and inaccuracies in the computer vision models. False negatives (i.e., the software failed to detect a completed movement) led to unnecessary delays, whereas false positives (i.e., incomplete movements incorrectly judged as complete) caused the software to advance prematurely.
Therapists noted that such pacing discrepancies, whether too slow or too fast, could lead to unnecessary exertion or discomfort. For example, if the software delays progression while verifying completion of an instruction such as “hold an object in the air,” a patient may maintain the position longer than clinically intended, increasing the risk of fatigue or muscle strain.
@@ 3 Discussion
We propose a novel digital intervention paradigm in which clinicians use LLMs to translate their exercise prescriptions into intervention software that can remotely instruct and monitor patients’ exercises. In our study, the generated software was executed on a standardized patient’s device to provide step-by-step guidance, monitor performance, and return monitored outcomes to clinicians for review. Our findings demonstrate that clinician-driven intervention software generation with LLM is feasible and acceptable to therapists in a standardized-patient setting, motivating larger trials to assess clinical effectiveness and safety in real-patient populations.
@@
The proposed digital intervention paradigm enabled a substantially larger fraction of personalized prescriptions to be delivered as software than the existing template-based paradigm, with a 45% increase in the proportion of prescribed exercises that could be implemented (100% vs. 55%; p <0.01). This improvement is an expected consequence of how personalization occurs in each paradigm. In our paradigm, therapists author software after the clinical encounter, once they have full knowledge of the patient’s impairments, environment, and goals, and use the LLM to generate code that directly reflects their prescriptions. By contrast, in the status quo, intervention software is implemented before the encounter, and therapists must retrofit their prescriptions into a fixed library of software, typically by selecting pre-existing software templates and adjusting a narrow set of parameters (for example, repetitions or duration) while the underlying exercise instructions remain fixed. As a result, many clinically important nuances—such as patient-specific household objects used in activities of daily living or contingency instructions when patients struggle with particular motions—cannot be represented, making the observed 45% gain in deliverable personalization both substantial and anticipated.
@@
Within this paradigm, the LLM-generated intervention software achieved 99.7% accuracy in delivering exercise instructions (397/398 instructions delivered as prescribed). This near-perfect accuracy is expected: the LLM is not designing exercises or making clinical decisions, but simply transcribing therapists’ prescriptions into executable software. As in standard practice, therapists reviewed the patient chart (here, a clinical vignette summarizing physical deficits in Sec. 4.2), interacted with the standardized patient to elicit needs, and prescribed personalized exercises. The LLM then functioned solely as a constrained translator. As specified in the system prompt (Fig.1), it was instructed to deliver exercise instructions verbatim as prescribed, allowing only minor clarifications (for example, resolving ambiguous pronouns such as “this” or “that” based on prior context). In this setting, the task primarily leverages the LLM’s strength in converting text descriptions into structured logic for software execution[llm_code_gen], rather than requiring complex reasoning. Our use of LLMs is therefore distinct from prior applications that rely on richer clinical reasoning—such as diagnosis[liu2025generalist], treatment planning [sandmann2025deepseek], or risk management [pais2024medic]. Therapists’ feedback further indicated that this constrained, code-generation role of LLM could be easily integrated into their existing clinical workflow and, in some cases, perceived as enhancing productivity.
@@
In contrast, monitoring whether patients follow exercise instructions correctly is substantially more challenging. It requires the system to interpret each instruction’s intent and relate it to patient’s movement data from computer vision models to decide whether the instruction has been completed. Implementing this monitoring logic demands both semantic understanding of the prescribed action and robust handling of noisy, real-world movement data, making it a harder problem than textual transcription. In this setting, we observed monitoring inaccuracies arising from hallucinations in the generated logic and limitations of the computer vision models. Existing DHIs typically rely on libraries that are manually implemented and tested by software engineers, which may, in principle, support higher monitoring reliability than automatically generated programs; our current implementation did not include a manual validation step. Despite these challenges, we observed an 88.4% monitoring accuracy (352/398 instructions; 95% CI, 0.84–0.92), which we view as a reasonable preliminary outcome that motivates further investigation of LLM-based software generation for digital interventions. The commercial LLM and computer vision models used here were not fine-tuned for this rehabilitation context; fine-tuning on domain-specific data and incorporating more robust pose-estimation and activity-recognition pipelines are likely to further improve monitoring performance as underlying models advance.
@@
Regarding patient safety, 90% (18/20) of therapists judged the LLM-generated intervention software to be safe in this standardized-patient setting. Because the safety of such software had not previously been evaluated, we deliberately used a standardized patient rather than real patients in this first feasibility study. This initial safety signal is encouraging but preliminary and requires confirmation in trials with real patients and longer-term use. Therapists observed the software articulate each prescribed instruction and wait until monitoring confirmed completion (or time-to-completion expired) before proceeding, enabling adaptive pacing aligned with patient performance and making sessions more interactive than reading step-by-step instructions on a screen. Most therapists considered the pacing adequate; however, two raised safety concerns about premature or delayed instruction transitions that could pose risks—for example, delayed detection of completing a task such as “hold an object in the air” could lead to fatigue or muscle strain. Improving monitoring accuracy through targeted fine-tuning may reduce premature transitions caused by false positives, while delayed transitions may be mitigated by selecting sensor-processing models—exposed to the LLM as APIs in the system prompt (Fig. 1)—with lower latency.
@@
To broadly support physical and occupational therapy with the proposed LLM-based paradigm, our study found that iterative co-design of the system prompt (Fig. 1) between software developers and clinicians was essential. The prompt specifies guidelines for the LLM to generate intervention software, detailing desired functionalities and guardrails aligned with clinical needs. Because authoring such prompts currently requires programming expertise, AI researchers or software engineers remain necessary contributors to developing this class of tools, as with other LLM-based clinical tools. In this study, however, the prompt only partially met therapist expectations, tempering adoption. Some therapists sought finer control over instruction pacing, while others expected monitoring beyond binary labels (complete vs. incomplete). Although 75% (15/20) of therapists indicated willingness to adopt the LLM software-generation tool, with unanimous consensus (20/20) on its ease of use, these unmet needs contributed to hesitancy, underscoring the importance of early needs-finding and iterative prompt refinement to ensure alignment with clinical practice.
@@
Our findings may have implications for other clinical domains that use DHIs. DHIs in mental health [personalization-ibis-review], chronic disease management [ai-diabetes-management], pain management [personalizing-pain-management], and speech–language pathology [digital-aphasia-therapy] face similar constraints on personalization. As an illustrative (untested) example, a study of exposure and response therapy for obsessive–compulsive disorder (OCD) [Lenhard2017] did not permit clinicians to encode new task-specific contingencies (for example, “proceed to Step 3 only if hands remain unwashed and visible for 10 minutes”) beyond predefined instructions. In contrast, our approach allows clinicians to specify such contingencies directly in the intervention software (Table 1). Prospective studies are needed to determine whether the benefits of LLM-based software generation observed here extend to other areas of healthcare.